An actor invests some money at 7​%, and ​$39000 more than twice the amount at 12%. The total annual interest earned from the investment is ​$32580. How much did he invest at each​ amount? Use the​ six-step method.
Answer by addingup(3677)

let "some money" be x:
0.07x + 0.12(2x+39000) = 32580
0.07x + 0.24x + 4680 = 32580
0.31x = 27900 = 90000 this is the amount invested at 7%
at 12%: 2(90000)+39000 = 219000